OpenStudy (anonymous):

HEEELPPP!!!!Give the equation of the circle with the radius 5 whose center is the the origin.

OpenStudy (richyw):

the equation for a circle centered at the point \((a,b)\) with radius \(r\) is\[(x-a)^2 + (y-b)^2 = r^2\]

OpenStudy (richyw):

and you know it is at the origin, so you have \(a,b\) and \(r\). Just plug it in!

OpenStudy (richyw):

Or in polar coordinates you can just say \(r(\theta)=5\)